# How Haag-tied is QFT, really?

### 18 Dec 2023 v2

@@ 1 Introduction
Set it up one last time, except now you are a philosopher of science. You may consider yourself a realist or an instrumentalist, interested in metaphysics or methodology—whatever comes to mind. Regardless, several things are true of you. First, you are committed to understanding the foundations of QFT (no matter what this means in practice). Second, you are familiar both with the major advances in conventional (Lagrangian) QFT that use the interaction picture as well as those based on algebraic QFT (AQFT). [^5] See [Fraser, 2011] and [Wallace, 2011] for the classic debate over these two formulations of QFT. But third, it is unclear to you if, or how, these advances can form a consistent whole.
@@
The remainder of the paper is structured as follows. In section 2 we provide a synopsis of a standard proof of Haag’s theorem and discuss the sense in which it raises an alarm that something is not right with the interaction picture. In section 3 we motivate the need for a framework for understanding the literature on Haag’s theorem. The framework itself is given in section 4. The framework employs and extends Hilbert’s construction analogy: the framework understands each author as something like a contractor giving their assessment of the problem in the foundations of physics heralded by Haag’s theorem, their recommendation for repair work on the foundations, and their expectations of the needed long-term maintenance or future renovations to QFT’s area of the edifice of science. This section applies the framework to seven leading contemporary viewpoints on Haag’s theorem; the key results of this application are given in table 1. Section 5 demonstrates the framework’s judicious balance of conceptual structure and flexibility in order to bring clarity to the space of responses to Haag’s theorem; it further argues that, at the end of the day, the most important lesson for philosophers to take from Haag’s theorem is that we need to put our own energies into clearly answering meta-level questions regarding the nature and purpose of foundational work in physics. Concluding remarks are given in section 6.
We will use the following terms to disambiguate the different approaches to and versions of QFT. By canonical QFT we mean what [Wallace, 2001] does by ‘Lagrangian QFT’: an approach that, for the most part, proceeds from specification of a classical field theory, to its quantization, and finally to its (Wilsonian) renormalization. By formal variants of QFT we mean the most familiar collection of approaches focused on mathematical rigor and precision, including algebraic QFT, constructive QFT (CQFT), and axiomatic QFT; these center around the Wightman, Osterwalder-Schrader, or Haag-Kastler axioms.
@@ 2.1 History of Haag’s Theorem
Haag’s theorem is the culmination of two approaches in early quantum field theory. On the one hand, Dyson had combined Feynman’s rules and the Tomonaga-Schwinger formalism into a reliable and practical approach for calculating the results of scattering experiments in relativistic quantum field theory. This approach brought with it a shift toward considering scattering amplitudes between free, asymptotic states (see [Blum, 2017]), there accomplished by the so-called interaction picture to model interactions [Schwinger, 1948b] (see 2.2). The approach was wildly successful, in particular in its use to calculate the anomalous magnetic moment of the electron [Schwinger, 1948a]. This is the genesis of canonical QFT.
@@
On the other hand, spurred by Wigner’s precise characterization of special relativity’s implications for quantum theory [Wigner, 1939], a more mathematical approach to understanding the structure of relativistic quantum field theories maintained an interest in determining the properties of instantaneous states in addition to free, asymptotic states. This is the genesis of formal variants QFT. Naturally, a transparent representation of relativistic QFT’s mathematical structure was prized, leading to an enumeration of several assumptions. However, the assumptions underlying the former approach’s use of the interaction picture were unclear. Haag, in his lecture series given at CERN in 1952–3, [^6] See [Haag, 2010] for his recollections of this aspect of the history. Note that Haag says (incorrectly) that the lectures were given in 1953–4; see [Lupher, 2005]. began by stating assumptions for models of relativistic QFT for a single, fundamental field. These included, following Wigner’s suggestion, requiring an irreducible Hilbert space representation of the Poincaré group. Then, Haag used the mathematical tools at his disposal to characterize the interaction picture (e.g., the particle representation of field states, unitary intertwiner). The problem (Haag’s theorem [Haag, 1955]) was that the interaction picture, thus characterized, could not be used to represent non-trivial interactions. This result was clarified and generalized by [Hall and Wightman, 1957], to which most discussions of Haag’s theorem refer. Essentially, Haag, Hall, and Wightman showed that “to give different physical predictions, two theories of an interacting field which satisfies the canonical commutation relations must use inequivalent representations of the commutation relations” [Hall and Wightman, 1957, 2].
@@
We should make two historical remarks. First, Haag’s theorem was not unique in observing the existence and importance of unitarily inequivalent representations of the CCRs. [Reed and Simon, 1975, 329] suggest—and [Haag, 1955, 21] asserts—that von Neumann may have known of the existence of inequivalent representations by the time he wrote [von Neumann, 1938], and, von Neumann notwithstanding, the particle representation originating in [Dirac, 1927] was the only one used until the late 1940s. But by 1951, Friedrichs had characterized a new representation and demonstrated its general inequivalence with the particle representation. He moreover argued that “in a certain sense [such fields] do occur in nature [Friedrichs, 1953, 142;153], and, in fact, the representation allows one to treat problems previously-insoluble because of the so-called infrared catastrophe. Similarly, [Wightman and Schweber, 1955] produced inequivalent representations (using von Neumann’s techniques) and suggest their physical relevance. Garding and Wightman then published their classification of all representations of the CARs in [Garding and Wightman, 1954], noting along the way that unitary equivalence also depends on the “terminal behavior” of the involved unitary operators when the measure (partially) characterizing the representation is nondiscrete [Garding and Wightman, 1954, 621]. This line of thinking led ultimately to Theorem X.46 of [Reed and Simon, 1975, 233], which says that free, neutral scalar fields of inequivalent mass carry unitarily inequivalent irreducible representations. This theorem will be relevant in §4, where e.g., [Duncan, 2012, Klaczynski, 2016, Earman and Fraser, 2006], construe this theorem as the conceptual core of Haag’s theorem.
@@
Second, the early reception of Haag’s theorem emphasized its implications for a particle interpretation but neither reduced its significance to its interpretive implications nor viewed it as ultimately damning for particles. It will suffice to discuss Haag himself. While he emphasized its consequences for the “possibility of defining a theory[…] which describes the interaction processes of particles,” he nevertheless (a) notes in summary that relaxing the equal-time vanishing of the CCRs suffices to evade the theorem and (b) claims the assumptions he makes are not only compatible but even “have physical significance” in the lowest orders of a perturbation expansion [Haag, 1955, 36-7]. Haag actually emphasizes two escape paths taken by later commentators, namely that his use of strong convergence diverges from the typical convergence factor and that satisfactory physical descriptions do not require use of infinite-dimensional Hilbert spaces. The point for Haag, it seems, was to probe the limits of the conventional field theory scheme [Haag, 1955, 32].
@@ 2.2 The Interaction Picture
Haag’s theorem is often construed as a no-go theorem for the interaction picture. As noted, this picture has been widely used for the calculation of many physical quantities that have matched experimental results to a high degree of accuracy, for example the celebrated computation of the anomalous magnetic dipole moment of the electron by [Schwinger, 1948a], and is a mainstay of undergraduate and graduate textbooks and has facilitated the calculation of many physical quantities that have matched experimental results to a high degree of accuracy. 777 As one crude measure of how salient the interaction picture is for actual physics, a search for the exact phrase “interaction picture” in Physics Reviews conducted on Nov 4, 2021 returned 4,377 hits. The interaction picture forms the basis of calculations using time-dependent perturbation theory. Crucially, the Gell-Mann and Low theorem makes use of the interaction picture. The original paper [Gell-Mann and Low, 1951] presenting this formula has (as of Nov 4, 2021) 1544 citations recorded on google scholar. This is likely a rather low estimate of how widely the formula is in fact used since the formula is now at the level of common knowledge among physicists, and thus in many cases it might be used without citation. For instance, the Gell-Mann and Low theorem was used to rigorously derive the Bethe–Salpeter equation, which describes the bound states of a system of two fermions in a relativistic formalism [Salpeter and Bethe, 1951]. This equation was used in subsequent quantum electrodynamics calculations of the fine structure of Hydrogen-like atoms [Salpeter, 1952], Helium [Douglas and Kroll, 1974] and heavy atoms [Mohr et al., 1998] and contemporary precision calculations of the electron magnetic moment [Aoyama et al., 2015]. Likewise, the equations have been utilized in the calculation of nucleon-nucleon potentials [Partovi and Lomon, 1969, Partovi and Lomon, 1972, Lomon, 1976, Partovi, 1978] and the derivation of scaling laws for interactions with large momentum transfer, confirmed by scattering experiments [Brodsky and Farrar, 1975], among many other examples. As its name suggests, the interaction picture is one way to model interacting fields in conventional quantum field theory. Let us suppose we have a field, $\phi$, with conjugate momentum $\pi$, generally taken (either explicitly or implicitly) to obey the equal time canonical commutation relations and other axioms of QFT.
The interaction picture is intermediate between the Schrödinger and Heisenberg pictures. In the Schrödinger picture of quantum mechanics, states evolve in time under the full Hamiltonian, whilst operators are stationary. In the Heisenberg picture of quantum mechanics, the operators evolve under the full Hamiltonian, whilst states are stationary. To form the interaction picture, we split the full Hamiltonian into a free and a (time-dependent) interaction part, $H=H_{F}+H_{I}$. The evolution of operators is governed by the free Hamiltonian, $H_{F}$, so the fields are free. The evolution of states is governed by the interaction part, $H_{I}$.
@@ 2.3 Proof of Haag’s Theorem for Spin-free, Neutral, Scalar Fields
The core strategy is due to [Hall and Wightman, 1957] rather than [Haag, 1955]. The main technical advance of the former over the latter was a series of theorems concerning the analytic continuation of functions invariant under the orthochronous inhomogeneous Lorentz group, which allowed them to extend the equality of two-, three-, and four-point vacuum expectation values at equal times to any times. For simplicity, we eschew reference to this complex, noting its intuitive gloss as “a quantitative formulation of the intuitive feeling that in a Lorentz invariant theory the equivalence of descriptions in different Lorentz frames should somehow restrict the possible correlations between the values of physical quantities at different points in space-time” [Hall and Wightman, 1957, 35]. We likewise appeal to the simpler Jost-Schroer theorem, which at any rate affords extension to all vacuum expectation values (when one of the fields is free). For convenience, we also adopt some of the notation and conventions of [Earman and Fraser, 2006] and [Seidewitz, 2017], including the numbering due to the latter.
@@ 2.3.1 Wightman Axioms
  2. 2.
    There is a unique, invariant vaccum state, |$\ket{0}$, in $\mathcal{H}$, invariant under $U$: $U(\mathbf{\Delta x},\Lambda)\ket{0}=\ket{0}$.
  3. 3.
@@ 2.3.2 Proof of Haag’s theorem
Haag’s theorem follows from the results of two other theorems. For clarity of exposition, here we only prove the theorem for neutral, scalar fields; the generalization for other types of fields follows in a straightforward manner. From here on, it is convenient to decompose four-vectors as $x=(t,\mathbf{x})$, where the right hand side are the temporal and spatial components, respectively.
@@ 2.4 Implications of Haag’s Theorem
By the criteria used by particle physicists, the interaction picture has undoubtedly produced numerous notable successes (see note 7). However, the interaction picture is generally taken to depend upon all of the assumptions needed for Haag’s theorem, including Poincaré invariance and the existence of a unitary operator relating the two fields, that is, that the fields are unitarily equivalent. The apparent clash between the interaction picture and Haag’s theorem arises as follows. If the fields, $\phi_{F}$ and $\phi_{I}$, obey the usual axioms of QFT, and we have a unitary operator intertwining the two fields at even a single time (as in equations 9 and 10), and $\phi_{F}$ is free, then according to Haag’s theorem $\phi_{I}$ must also be free. So the interaction must be trivial.
@@ 3 What the Haag Is Going on?
Second, there is a consensus that Haag’s theorem poses a bona fide problem for the standard presentation of the interaction picture. [^10] That is, the pre-renormalization interaction picture. The theorem itself is mathematically correct: as Klaczynski puts it, “it is a mathematical theorem in the truest sense of the word; it brings with it the ‘hardness of the logical must’.” [Klaczynski, 2016]. Further, it is not disputed that the assumptions of Haag’s theorem hold in the case of the standard textbook presentations of the interaction picture. [^11] For example, standard presentations assume, either implicitly or explicitly, a Poincaré invariant theory, and that there are distinct free and interacting fields, related by a unitary operator. No one, to our knowledge, argues that the problem posed by Haag’s theorem is simply illusory. [^12] It may very well be that many practicing physicists would say that the problem, if not illusory, is trivial. Perhaps many physicists think that the problem posed by Haag’s theorem is at most rooted in an unrealistic idealization of non-interacting states at temporal infinity. And perhaps they therefore choose not to address Haag’s theorem in their textbooks, lecture notes, or research articles. We can do little more than speculate that such an unspoken consensus fully explains the dearth of references to Haag’s theorem in standard accounts of QFT—simple ignorance of the theorem may be just as strong of a causal factor. David Tong’s lecture notes on QFT, for instance, do not explicitly mention Haag’s theorem; but they do say that the assumption of non-interacting states in the interaction picture is wrong and should be replaced by the interactions-are-always-on interpretation of the LSZ reduction formula [Tong, 2012] p. 54-55, 79-80. We might tentatively read this as an attitude towards Haag’s theorem as minimally useful: it points out that an initial assumption was wrong, and in wise pedagogical fashion, Tong will correct that assumption in due course when more complexity and sophistication can be digested. But no need to belabor the point by naming and deeply discussing the theorem. Thus, while it is possible that this sort of largely dismissive response to Haag’s theorem is widespread, too little of it exists in print to be extensively covered in the remainder of this paper. It is, rather, the severity and appropriate remedy of the problem that is subject to debate.
@@
However, the severity of the problem raised by Haag’s theorem clearly stops short of spelling the demise of that research program predicated on the application of quantum theories of fields to scattering experiments, commonly called particle physics. [^13] Though see the discussion of Kastner (section 4.6) and Seidewitz (section 4.7) below for genuine proposals of new physical theories, each drawing some motivation from Haag’s theorem. Even the practitioners of AQFT, while operating far afield from the details of the Standard Model’s phenomenology, still conceive of their work as contributing to the scientific enterprise whose primary goal is to develop the quantum theory of fields as the appropriate theoretical apparatus for understanding scattering experiments. So, then, what is going on with Haag’s theorem such that these two stances —the interaction picture has been used successfully, and Haag’s theorem poses a substantive problem for the interaction picture—can be held together? This central tension is widely recognized as a point of agreement. As Teller succinctly puts it,
@@
The reader is thus confronted with paths out of the Haagian labyrinth diverging on both metaphysical and mathematical grounds and, worse still, she can’t tell from her place in the labyrinth whether or where these paths coincide. As if this weren’t bad enough, a fog sets in. Are we even trapped at all?, the reader asks, [Seidewitz, 2017, 356] in hand, for Haag’s theorem arises in traditional QFT only because time is not “treated comparably to the three space coordinates, rather than as an evolution parameter.” Thus, the Haagian labyrinth could have been entirely avoided had time been treated in a relativistically sensible manner at the outset.

### 06 Apr 2024 v3

@@ 1 Introduction
Set it up one last time, except now you are a philosopher of science. You may consider yourself a realist or an instrumentalist, interested in metaphysics or methodology—whatever comes to mind. Regardless, several things are true of you. First, you are committed to understanding the foundations of QFT (no matter what this means in practice). Second, you are familiar both with the major advances in canonical (Lagrangian) QFT that use the interaction picture as well as those based on algebraic QFT (AQFT). [^4] See [Fraser, 2011] and [Wallace, 2011] for the classic debate over these two formulations of QFT. But third, it is unclear to you if, or how, these advances can form a consistent whole.
@@
The remainder of the paper is structured as follows. In section 2 we offer a cursory historical overview, a synopsis of a standard proof of Haag’s theorem, and discuss the sense in which it raises an alarm that something is not right with the interaction picture. In section 3 we motivate the need for a framework for understanding the literature on Haag’s theorem. The framework itself is given in section 4. The framework employs and extends Hilbert’s construction analogy: the framework understands each author as something like a contractor giving their assessment of the problem in the foundations of physics heralded by Haag’s theorem, their recommendation for repair work on the foundations, and their expectations of the needed long-term maintenance or future renovations to QFT’s area of the edifice of science. This section applies the framework to seven leading contemporary viewpoints on Haag’s theorem; the key results of this application are given in table 1. Section 5 demonstrates the framework’s judicious balance of conceptual structure and flexibility in order to bring clarity to the space of responses to Haag’s theorem; it further argues that, at the end of the day, the most important lesson for philosophers to take from Haag’s theorem is that we need to put our own energies into clearly answering meta-level questions regarding the nature and purpose of foundational work in physics. Concluding remarks are given in section 6.
We will use the following terms to disambiguate the different approaches to and versions of QFT. By canonical QFT we mean what [Wallace, 2001] does by Lagrangian QFT: an approach that, for the most part, proceeds from specification of a classical field theory, to its quantization, and finally to its (Wilsonian) renormalization. By formal variants of QFT we mean the most familiar collection of approaches focused on mathematical rigor and precision, including algebraic QFT, constructive QFT (CQFT), and axiomatic QFT; these center around the Wightman, Osterwalder-Schrader, or Haag-Kastler axioms.
@@ 2.1 History of Haag’s Theorem
Haag’s theorem is the culmination of two approaches in early quantum field theory. On the one hand, Dyson had combined Feynman’s rules and the Tomonaga-Schwinger formalism into a reliable and practical approach for calculating the results of scattering experiments in relativistic quantum field theory. This approach brought with it a shift toward considering scattering amplitudes between free, asymptotic states (see [Blum, 2017]), there accomplished by the so-called interaction picture to model interactions [Schwinger, 1948b] (see 2.2). The approach was wildly successful, in particular in its use in calculating the anomalous magnetic moment of the electron [Schwinger, 1948a]. This is the genesis of canonical QFT.
@@
On the other hand, spurred by Wigner’s precise characterization of special relativity’s implications for quantum theory [Wigner, 1939], a more mathematically rigorous approach to understanding the structure of relativistic quantum field theories maintained an interest in determining the properties of instantaneous states in addition to free, asymptotic states. This was the genesis of formal variants QFT. Naturally, a transparent representation of relativistic QFT’s mathematical structure was prized, leading to an enumeration of several assumptions. However, the assumptions underlying the former approach’s use of the interaction picture were unclear. Haag, in his lecture series given at CERN in 1952–3, [^5] See [Haag, 2010] for his recollections of this aspect of the history. Note that Haag says (incorrectly) that the lectures were given in 1953–4; see [Lupher, 2005]. began by stating assumptions for models of relativistic QFT for a single, fundamental field. These included, following Wigner’s suggestion, requiring an irreducible Hilbert space representation of the Poincaré group. Then, Haag used the mathematical tools at his disposal to characterize the interaction picture (e.g., the particle representation of field states, unitary intertwiner). The problem (Haag’s theorem [Haag, 1955]) was that the interaction picture, thus characterized, could not be used to represent non-trivial interactions. This result was clarified and generalized by [Hall and Wightman, 1957], to which most discussions of Haag’s theorem refer. Essentially, Haag, Hall, and Wightman showed that “to give different physical predictions, two theories of an interacting field which satisfies the canonical commutation relations must use inequivalent representations of the commutation relations” [Hall and Wightman, 1957, 2].
@@
We should make two historical remarks. First, Haag’s theorem was not unique in observing the existence and importance of unitarily inequivalent representations of the CCRs. [Reed and Simon, 1975, 329] suggest—and [Haag, 1955, 21] asserts—that von Neumann may have known of the existence of inequivalent representations by the time he wrote [von Neumann, 1938], and, von Neumann notwithstanding, the particle representation originating in [Dirac, 1927] was the only one used until the late 1940s. But by 1951, Friedrichs had characterized a new representation and demonstrated its general inequivalence with the particle representation. He moreover argued that “in a certain sense [such fields] do occur in nature [Friedrichs, 1953, 142;153], and, in fact, the representation allows one to treat problems previously insoluble because of the so-called infrared catastrophe. Similarly, [Wightman and Schweber, 1955] produced inequivalent representations (using von Neumann’s techniques) and suggest their physical relevance; this built, in part, on a model van Hove used to demonstrate that interacting and free state spaces can be orthogonal [Van Hove, 1952]. [^6] See §7.2.2 of [Sbisà, 2020] for an explanation of how van Hove’s result complements, but does not follow from, Haag’s theorem. We also note with Sbisà that van Hove’s result implicates ultraviolet divergences, unlike those of Friedrichs or Haag (as Friedrichs himself noted [Friedrichs, 1953, 270-72]). Garding and Wightman then published their classification of all representations of the CARs in [Garding and Wightman, 1954], noting along the way that unitary equivalence also depends on the “terminal behavior” of the involved unitary operators when the measure (partially) characterizing the representation is nondiscrete [Garding and Wightman, 1954, 621]. This line of thinking led to results like Theorem X.46 of [Reed and Simon, 1975, 233], which says that free, neutral scalar fields of inequivalent mass carry unitarily inequivalent irreducible representations. This theorem will be relevant in §4, where e.g., [Duncan, 2012, Klaczynski, 2016, Earman and Fraser, 2006], construe this theorem as the conceptual core of Haag’s theorem.
@@
Second, the early reception of Haag’s theorem focused on its implications for a particle interpretation but neither reduced its significance to its interpretive implications nor viewed it as ultimately damning for particles. It will suffice to discuss Haag himself. While he emphasized its consequences for the “possibility of defining a theory[…] which describes the interaction processes of particles,” he nevertheless (a) notes in summary that relaxing the equal time vanishing of the CCRs suffices to evade the theorem and (b) claims the assumptions he makes are not only compatible but even “have physical significance” in the lowest orders of a perturbation expansion [Haag, 1955, 36-7]. Haag actually emphasizes two escape paths taken by later commentators, namely that his use of strong convergence diverges from the typical convergence factor and that satisfactory physical descriptions do not require use of infinite dimensional Hilbert spaces. The point for Haag, it seems, was to probe the limits of the conventional field theory scheme [Haag, 1955, 32].
@@ 2.2 The Interaction Picture
Haag’s theorem is often construed as a no-go theorem for the interaction picture. As previously noted, the interaction picture is a mainstay of undergraduate and graduate textbooks, and has facilitated the calculation of many physical quantities that have matched experimental results to a high degree of accuracy, 777, for example the celebrated computation of the anomalous magnetic dipole moment of the electron by [Schwinger, 1948a]. As one crude measure of how salient the interaction picture is for actual physics, a search for the exact phrase “interaction picture” in Physics Reviews conducted on Nov 4, 2021 returned 4,377 hits. The interaction picture forms the basis of calculations using time-dependent perturbation theory. Crucially, the Gell-Mann and Low theorem makes use of the interaction picture. The original paper [Gell-Mann and Low, 1951] presenting this formula has (as of February 4, 2024) 1637 citations recorded on google scholar. This is likely a rather low estimate of how widely the formula is in fact used since the formula is now at the level of common knowledge among physicists, and thus in many cases it might be used without citation. For instance, the Gell-Mann and Low theorem was used to rigorously derive the Bethe–Salpeter equation, which describes the bound states of a system of two fermions in a relativistic formalism [Salpeter and Bethe, 1951]. This equation was used in subsequent quantum electrodynamics calculations of the fine structure of Hydrogen-like atoms [Salpeter, 1952], Helium [Douglas and Kroll, 1974] and heavy atoms [Mohr et al., 1998] and contemporary precision calculations of the electron magnetic moment [Aoyama et al., 2015]. Likewise, the equations have been utilized in the calculation of nucleon-nucleon potentials [Partovi and Lomon, 1969, Partovi and Lomon, 1972, Lomon, 1976, Partovi, 1978] and the derivation of scaling laws for interactions with large momentum transfer, confirmed by scattering experiments [Brodsky and Farrar, 1975], among many other examples.
As its name suggests, the interaction picture is one way to model interacting fields in canonical quantum field theory. In the Schrödinger picture of quantum mechanics, states evolve in time under the full Hamiltonian, whilst operators are stationary. In the Heisenberg picture of quantum mechanics, the operators evolve under the full Hamiltonian, whilst states are stationary. The interaction picture is intermediate between these two pictures. To form the interaction picture, we split the full Hamiltonian into a free and a (time-dependent) interaction part, $H=H_{F}+H_{I}$. The evolution of operators is governed by the free Hamiltonian, $H_{F}$, so the fields are free. The evolution of states is governed by the interaction part, $H_{I}$.
@@ 2.3 Proof of Haag’s Theorem for Spin-free, Neutral, Scalar Fields
Haag’s theorem follows from the results of two other theorems. For clarity of exposition, here we only prove the theorem for neutral, scalar fields; the generalization for other types of fields follows in a straightforward manner. The core strategy is due to [Hall and Wightman, 1957] rather than [Haag, 1955]. The main technical advance of the former over the latter was a series of theorems concerning the analytic continuation of functions invariant under the orthochronous inhomogeneous Lorentz group, which allowed them to extend the equality of two-, three-, and four-point vacuum expectation values at equal times to any times. For simplicity, we eschew reference to this complex, noting its intuitive gloss as “a quantitative formulation of the intuitive feeling that in a Lorentz-invariant theory the equivalence of descriptions in different Lorentz frames should somehow restrict the possible correlations between the values of physical quantities at different points in space-time” [Hall and Wightman, 1957, 35]. We likewise appeal to the simpler Jost-Schroer theorem, which at any rate affords extension to all vacuum expectation values (when one of the fields is free). For convenience, we also adopt some of the notation and conventions of [Earman and Fraser, 2006] and [Seidewitz, 2017], including the numbering due to the latter.
@@ 2.3.1 Wightman Axioms
  2. 0.2
    There is a unique, invariant vaccuum state, |$\ket{0}$, in $\mathcal{H}$, invariant under $U$: $U(\mathbf{\Delta x},\Lambda)\ket{0}=\ket{0}$.
  3. 0.3
@@ 2.3.2 Proof of Haag’s theorem
From here on, it is convenient to decompose four-vectors as $x=(t,\mathbf{x})$, where the right hand side are the temporal and spatial components, respectively.
@@ 2.4 Implications of Haag’s Theorem
By the criteria used by particle physicists, the interaction picture has undoubtedly produced numerous notable successes (see note 7). However, the interaction picture is generally taken to depend upon all of the assumptions needed for Haag’s theorem, including Poincaré invariance and the existence of a unitary operator relating the two fields, that is, that the fields are unitarily equivalent. The apparent clash between the interaction picture and Haag’s theorem arises as follows. If the fields, $\phi_{F}$ and $\phi_{I}$, obey the Wightman axioms of QFT, and we have a unitary operator intertwining the two fields at even a single time (as in equations 9 and 10), and $\phi_{F}$ is free, then according to Haag’s theorem $\phi_{I}$ must also be free. So the interaction must be trivial.
@@ 3 What the Haag Is Going on?
Second, there is a consensus that Haag’s theorem poses a bona fide problem for the standard presentation of the interaction picture. [^10] More precisely, it is agreed to pose a problem for the pre-renormalization interaction picture. The theorem itself is mathematically correct: as Klaczynski puts it, “it is a mathematical theorem in the truest sense of the word; it brings with it the ‘hardness of the logical must’.” [Klaczynski, 2016]. Further, it is not disputed that the assumptions of Haag’s theorem hold in the case of the standard textbook presentations of the interaction picture. [^11] For example, standard presentations assume, either implicitly or explicitly, a Poincaré-invariant theory, and that there are distinct free and interacting fields, related by a unitary operator. No one, to our knowledge, argues that the problem posed by Haag’s theorem is simply illusory. [^12] It may very well be that many practicing physicists today would say that the problem, if not illusory, is inconsequential. Perhaps many physicists think that the problem posed by Haag’s theorem is at most rooted in an unrealistic idealization of non-interacting states at temporal infinity. And perhaps they therefore choose not to address Haag’s theorem in their textbooks, lecture notes, or research articles. Although several textbooks from the 1960’s give brief assessments of the theorem, more recent treatments rarely mention it (see [Earman and Fraser, 2006, 307-309]). We can do little more than speculate that such an unspoken consensus fully explains the dearth of references to Haag’s theorem in standard accounts of QFT—simple ignorance of the theorem may be just as strong of a causal factor. David Tong’s lecture notes on QFT, for instance, do not explicitly mention Haag’s theorem; but they do say that the assumption of non-interacting states in the interaction picture is wrong and should be replaced by the interactions-always-on interpretation of the LSZ reduction formula [Tong, 2012, 54-55, 79-80]. We might tentatively read this as an attitude towards Haag’s theorem as minimally useful: it points out that an initial assumption was wrong, and in wise pedagogical fashion, Tong will correct that assumption in due course when more complexity and sophistication can be digested. But no need to belabor the point by naming and deeply discussing the theorem. Thus, while it is possible that this sort of largely dismissive response to Haag’s theorem is widespread, too little of it exists in print to be extensively covered in the remainder of this paper. It is, rather, the severity and appropriate remedy of the problem that is subject to debate.
@@
However, the severity of the problem raised by Haag’s theorem clearly stops short of spelling the demise of that research program predicated on the application of quantum theories of fields to scattering experiments, commonly called particle physics. [^13] Though see the discussion of Kastner (section 4.6) and Seidewitz (section 4.7) below for genuine proposals of new physical theories, each drawing some motivation from Haag’s theorem. Even the practitioners of AQFT, operating far afield from the details of the Standard Model’s phenomenology, still conceive of their work as contributing to the scientific enterprise whose primary goal is to develop the quantum theory of fields as the appropriate theoretical apparatus for understanding scattering experiments. So, then, what is going on with Haag’s theorem such that these two stances —the interaction picture has been used successfully, and Haag’s theorem poses a substantive problem for the interaction picture—can be held together? This central tension is widely recognized as a point of agreement. As Teller succinctly puts it,
@@
The reader is thus confronted with paths out of the Haagian labyrinth diverging on both metaphysical and mathematical grounds and, worse still, she cannot tell from her place in the labyrinth whether or where these paths coincide. As if this were not bad enough, a fog sets in. Are we even trapped at all?, the reader asks, [Seidewitz, 2017, 356] in hand, for Haag’s theorem arises in traditional QFT only because time is not “treated comparably to the three space coordinates, rather than as an evolution parameter.” Thus, the Haagian labyrinth could have been entirely avoided had time been treated in a relativistically sensible manner at the outset.